Suppose you take a 30-year mortgage of $300000. the annual interest rate is 4%, and the annual ... Question: Suppose you take a 30-year mortgage of $300000. The annual interest rate is 4%, and the annual AP... (2 bookmarks) Suppose you take a 30-year mortgage of $300000. The annual interest rate is 4%, and the annual APR is 5.00%. Loan payments are made annually. Calculate the amortized fees and expenses for this loan (in dollars, provide your answer with $1 precision). Answer: 2166 +/- 1

This is the steps to solve this problem, but I want to know how to do it by hand or calculator by hand(not excel) Please help!

Given loan term=30 years =nper

Loan amount =$300,000=PV

Annual interest rate=4%.=rate

So, the periodic payment can be calculated using PMT function of excel=PMT(rate,nper,pv,[Fv],[type])

Substituting the values in the formula

=PMT(.04,30,-300000,,0) HOW TO DO THIS<-----

=$17,349.03.---------(1)

If the APR is 5%, then effective cost of the loan is calculated by replacing 4% with 5% in the above formula ie.,

=PMT(.05,30,-300000,,0) AND HOW TO DO THIS<-----

=$19,515.43.----------(2)

The difference between (2) and (1) is nothing but the amortized fees and expenses i.e,

$19,515.43 - $17,349.03.

=$2,166.40
Answer #1

1:

Loan term=30 years =N

Loan amount =$300,000=PV

Annual interest rate=4%.=r

PV of annuity = A*(1-1/(1+rate)^number of terms)/rate

300000 = A*(1-1/1.04^30)/0.04

300000 = A*17.29203

A = 17349.03

2:

At rate = 5%

PV of annuity = A*(1-1/(1+rate)^number of terms)/rate

300000 = A*(1-1/1.05^30)/0.05

300000 = A* 15.37245

A = 19515.43